OpenStudy (anonymous):

solve the following equation e^1-8x=7957

OpenStudy (lncognlto):

Solving for x from e^1-8x=7957 gives us: e^1 - 7957 = 8x. therefore x = (e^1 - 7957)/8 Working this out on a calculator gives us: x = -994.2852148...

OpenStudy (lncognlto):

\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)